OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5

OpenStudy (dan815):

you can multiply both be 2 or any constant and they will still be the same system

OpenStudy (anonymous):

A. -4x - 9y = -19 -10y = -30 B. 4x + 3y = 5 2y = -6 C.7x - 3y = -11 9x = -6 D. 12x - 3y = -33 14x = -28

OpenStudy (dan815):

you can also make a new equation which is the addition of both of these equations and that can be one new equation you can use

OpenStudy (dan815):

D. 12x - 3y = -33 14x = -28

OpenStudy (dan815):

the first equation got multiply by 3 and then for the 2nd equation u add the first equatio nand the 2nd equation,

OpenStudy (dan815):

do you want more of an explanation

OpenStudy (dan815):

( 4x - y = -11 ) *3 12x-3y=-33 -12x+3y=33 <-- multiply everything 1 now look at this -14x+2x+3y=28+5, all i did was rewrite hte -12 to -14+2 now look at this part in the middle we know that 2x+3y =5 already so -14x+(2x+3y)=28+(5) -14x=28 <--- this means this has to be true 2x + 3y = 5

OpenStudy (dan815):

3rd step should say multiply everything by -1*